# Qubit-Efficient Quantum Annealing for Stochastic Unit Commitment

### v1

@@ II-B Quantum Computing
With the maturity of quantum technology, the currently available quantum annealers have the potential to solve large-scale integer programming problems. This section provides conversion of classical QUBO problems into quantum forms, followed by the principles and challenges of QA algorithm.
@@ II-B1 QUBO Problem and Hamiltonian Representation
QUBO, an NP-hard problem, is one of the most widely studied optimization problems in QC. In this model, the solution to the optimization problem is composed of a series of binary variables whose value is 00 or $1$, and these variables correspond exactly to the off or on state of units in (2). The general expression of the QUBO problem is given as
@@
where =$\aleph=\{1,\ldots,\mathfrak{n}\}$ is a set of the binary variable indices. Then, $\tau_{i},\tau_{j}\in\{0,1\}$ are binary variables, and $B_{ij},c_{i}\in\mathbb{R}$ are the coefficients of quadratic and linear terms, respectively. To allow QUBO to be executable by quantum compter, it must be mapped into Ising model. The Ising-Lenz model is one of the prevalently used methods, which originally intended to explain the phase transition of ferromagnets in statistical mechanics [30]. In quantum mechanics, the binary variables in the Ising model represent the spin direction of particles. The spin-up state $\left|\uparrow\right\rangle=\begin{bmatrix}1&0\end{bmatrix}^{\mathrm{T}}$ and spin-down state $\left|\downarrow\right\rangle=\begin{bmatrix}0&1\end{bmatrix}^{\mathrm{T}}$ represent $\upsilon$ correspond to $+1$ and $-1$, respectively [11]. In QC, we define $\left|0\right\rangle$ is spin up and $\left|1\right\rangle$ is spin down. Hence, the substitution relationship between QUBO and the Ising model is $\tau=(1-\upsilon)/2$. The expression of the Ising model can then be derived from (4) as
@@
In QC, all results are obtained through measurement, in contrast to classical computing. Therefore, two key aspects of measurement require precise definitions.
The first is the representation. In a quantum system, the probability distribution characteristics of particles are clarified by wave functions, and the expansion of the wave functions varies depending on the chosen representation. Hence, useful measurement results only occurs when the representation is clearly defined. Generally, the spin direction along the z-axis is chosen as the representation in QC.
The next necessary concept is the measurement operator. According to the definition in quantum mechanics, any measurable physical quantity can be characterized by a Hermitian operator, which is a square matrix acting on the state space. Additionally, once the measurement operator is determined, all possible measurement results must be eigenvalues of the observable. From this, it is evident that the key to selecting the measurement operator lies in matching the appropriate eigenstates and eigenvalues. As mentioned above, spin up $\left|\uparrow\right\rangle=\begin{bmatrix}1&0\end{bmatrix}^{\mathrm{T}}$ and spin down $\left|\downarrow\right\rangle=\begin{bmatrix}0&1\end{bmatrix}^{\mathrm{T}}$ are eigenstates in the Ising model, with corresponding eigenvalues of 1 and -1. Thus, we can use Pauli-z operator $\boldsymbol{\sigma}^{z}$ to explicitly delineate our measurement space. By replacing $\upsilon$ with $\boldsymbol{\sigma}^{z}$ in (5), the quantum expression of the QUBO objective function becomes a Hermitian operator called Hamiltonian, which is composed of Pauli-z operators as
$\boldsymbol{H}_{QUBO}$ $=\sum_{i\in\aleph}\sum_{j\in\aleph,i\neq j}Q_{ij}\boldsymbol{\sigma}_{i}^{z}\otimes\boldsymbol{\sigma}_{j}^{z}+\sum_{i\in\aleph}p_{i}\boldsymbol{\sigma}_{i}^{z},\;\text{with}$ (6)
$\boldsymbol{\sigma}^{z}$ $=\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right]$
II-B2 Quantum Annealing
QA utilizes the quantum tunneling effect, induced by quantum fluctuations, to overcome suboptimal local solutions, enabling the search for the global minimum of an objective function among a vast number of candidate local minimum solutions [13]. Quantum tunneling effect can tunnel through barriers, transitioning directly from a local minimum to the global minimum. Therefore, the QA algorithm usually outperforms classical computing in solving QUBO problems.
@@ II-B2 Quantum Annealing
QA design is grounded in the principle of adiabatic computation. This theory stresses that if the initial qubits are in the eigenstate of the initial Hamiltonian, the final quantum state will evolve into the ground state of the objective Hamiltonian when the evolution of quantum systems is sufficiently slow and no energy level crossings occur [18, 31]. When the objective Hamiltonian corresponds to the Ising model mapped from the objective function, the final ground state of QA can be regarded as the global optimum of the objective function. However, if the barriers in the problem are too wide, it is possible for the system to evolve into a local optimum. The mathematical expression of this process is as
$\boldsymbol{H}_{mixed}(t)=[1-s(t)]\boldsymbol{H}_{I}+s(t)\boldsymbol{H}_{F}$ (7)
@@
where $\boldsymbol{H}_{I}$ is an initial Hamiltonian, $\boldsymbol{H}_{F}$ is a final Hamiltonian that equals the objective Hamiltonian, and $s(t)$ is referred to as the normalized evolution time.
The described QA procedure can currently be executed on D-Wave QA processors. The processing units of the D-Wave 2000Q quantum computer consist of lattices made up of tiny metal loops, with each loop representing a quantum bit. The interconnection structure between qubits is known as the Chimera graph, which is composed of interconnected unit cells, with each unit cell containing 8 qubits. Once the QUBO expression of the objective function is formulated, the corresponding graph embedding can be constructed on the D-Wave QA-QPU, enabling the identification of the eigenstate with the lowest energy.
II-B3 Challenges of QA
Despite D-Wave QA-QPUs, such as the Advantage_system 4.1, offering a substantial qubit count (up to 5,627 qubits) [32], practical applications are still limited. First, the non-fully connected topology of QA-QPUs requires more qubits for embedding than the actual QUBO variables in theory, as additional qubits are needed to map the objective function correctly. For example, achieving a fully connected topology for 4 binary variables in a QUBO problem may require 6 qubits in the QA-QPU as Fig. 1 showing. Second, QA is affected by the chain-breaking issue [33], where the embedding process fails to preserve the integrity of the QUBO representation, leading to erroneous annealing outcomes. These limitations significantly restrict the scalability and effectiveness of QA in directly solving large-scale optimization problems.
Figure 1: 4-qubit full-connected QUBO problem and the embedded structure in Chimera unit cell (Ignore the chain strength embedding).
@@ III Proposed QPHR-ADMM Algorithm
As highlighted in the previous section, designing a QA algorithm with minimal qubit overhead is crucial for addressing real-world problems. This section first introduces binary encoding and augmented Lagrangian function to transform (2) into QUBO form and then demonstrates that a direct implementation of QA for solving (2) inevitably requires an increasing number of additional qubits as the Benders’ iterations progress, compounding the already significant qubit demand introduced by SUC itself. Two novel algorithms, namely PHR Augmented Lagrangian and quantum-based ADMM, are proposed to resolve the challenges with theoretical analysis on their space complexity.
III-A Pre-processing for Quantumization
As the master problem in Section II-A is a constrained MIP problem rather than in a QUBO form, QA-QPUs cannot directly construct the corresponding quantum embedding. Therefore, preprocessing techniques are presented in this section.
@@ III-A1 Binary Encoding
To start, the continuous variable $\Upsilon$, introduced by Benders LB (2), cannot be directly processed by the QPU. Binary encoding [34], expressed as $\Upsilon^{k}=\chi\sum_{j=0}^{J-1}2^{j}u_{j}^{k}$, can be applied where $u_{j}^{k}$ are binary variables. By selecting the appropriate precision coefficients $\chi$ and precision levels $J$, the desired precision can be attained. Leveraging the $NT$ number of binary variables in the original SUC, the qubit occupation after binary encoding becomes $NT+J$.
@@ III-A2 Augmented Lagrangian Function
To eliminate the inequality Bender’s cut (2c), augmented Lagrangian can be used by introducing extra slack variables into the objective. However, since slack variables are continuous, they also require binary encoding. According to (2c), the number of Benders cuts grows with the number of iterations $k$. Assuming the precision level of each slack variable is $\mathcal{F}$, the qubit overhead for the QPU grows linearly as iteration $k$ increase and can be denoted as $NT+J+k\mathcal{F}$.
Considering the limited number of qubits in quantum computers, introducing slack variables is inefficient and unacceptable. This inefficiency serves as the primary motivation for the PHR Lagrangian algorithm proposed in this Section III-B. Moreover, the qubit occupation grows linearly with the number of generator $N$. This qubit overhead will be resolved by the quantum ADMM algorithm in III-C.
@@ III-B1 Classic Formulation
The Powell-Hestenes-Rockafellar Augmented Lagrangian Multiplier (PHR-ALM) method was originally designed for equality-constrained optimization problems. Rockfellar later extended it to handle inequality constraints, and this method eliminates the need for slack variables in optimization problems [35], which is an ideal property for QA.
@@
$\min$ $f(\boldsymbol{x})$ (8)
s.t. $g_{i}(\boldsymbol{x})\leq 0\text{ for }i=1,\ldots,w$
where $n$ represents the total number of unit states and binary-encoded variables. $\boldsymbol{x}\in\{0,1\}^{n}$ and $w$ is the total number of binary inequality constraints after $w$ BD iterations. After introducing slack variables $\mathcal{S}_{i}$ to relax each inequality constraints, the augmented Lagrangian function $\mathcal{L}$ of (8) can be given as
$\mathcal{L}$ $=f(\boldsymbol{x})\!+\kappa(\boldsymbol{x},\mathcal{S},\lambda,\sigma)$ (9)
$=f(\boldsymbol{x})\!+\lambda_{i}\sum_{i=1}^{w}\left[g_{i}(\boldsymbol{x})\!+\mathcal{S}_{i}{}^{2}\right]\!+\frac{\sigma}{2}\sum_{i=1}^{w}\left[g_{i}(\boldsymbol{x})\!+\mathcal{S}_{i}{}^{2}\right]^{2}$
@@
$\nabla_{\mathcal{S}}\mathcal{L}=2\lambda_{i}\mathcal{S}_{i}+2\sigma\mathcal{S}_{i}\left[g_{i}(\boldsymbol{x})+\mathcal{S}_{i}{}^{2}\right]=0\quad\forall i$ (10)
@@
${\mathcal{S}_{i}{}^{2}}=\begin{cases}\hfil{-\frac{\lambda_{i}}{\sigma}-g_{i}(\boldsymbol{x})},&\lambda_{i}+\sigma g_{i}(\boldsymbol{x})\leq 0\\ \hfil{0,}&{\lambda_{i}+\sigma g_{i}(\boldsymbol{x})>0}\end{cases}$ (11)
From (11), when $\lambda_{i}+\sigma g_{i}(\boldsymbol{x})\leq 0$, $\mathcal{S}_{i}{}^{2}$ is non-negative, which means that the solution satisfies the $i$-th constraint $g_{i}(\boldsymbol{x})\leq 0$. In contrast, it indicates that the solution is not within the feasible region. Next, substituting the above results into the augmented penalty function, the corresponding PHR-based expression is given as follows,
${\kappa_{i}}=\begin{cases}\hfil-\frac{\lambda_{i}{}^{2}}{2\sigma},&\lambda_{i}+\sigma g_{i}(\boldsymbol{x})\leq 0\\ \hfil{\frac{1}{2\sigma}\left[\left(\sigma g_{i}(\boldsymbol{x})+\lambda_{i}\right)^{2}-\lambda_{i}{}^{2}\right],}&{\lambda_{i}+\sigma g_{i}(\boldsymbol{x})>0}\end{cases}$ (12)
@@
$\mathcal{L}=f(\boldsymbol{x})+\frac{1}{2\sigma}\sum_{i=1}^{w}\left(\left[\max\left\{\sigma g_{i}(\boldsymbol{x})+\lambda_{i},0\right\}\right]^{2}-\lambda_{i}{}^{2}\right)$ (13)
@@
At iteration $\ell$, solving the unconstrained PHR augmented Lagrangian function (13) yields the solution $\boldsymbol{x}^{\ell}$. If $\boldsymbol{x}^{\ell}$ satisfies the stop criterion (14c), the iteration can be terminated, and the output $\boldsymbol{x}^{\ell}$ is the approximate minimum solution of the original problem. Otherwise, update the next step Lagrange multipliers $\lambda_{i}^{\ell+1}$ and the penalty coefficient $\sigma^{\ell+1}$ according to the rules in (14a) and (14b) for the next iteration.
$\lambda_{i}^{\ell+1}$ $=\max\!\left\{\lambda_{i}^{\ell}\!+\sigma^{\ell}g_{i}\left(\boldsymbol{x}^{\ell+1}\right)\!,0\right\}\text{ for }i=1,\ldots,w$ (14a)
$\sigma^{\ell+1}$ $=\eta\sigma^{\ell}$ (14b)
$\mathcal{R}^{\ell}$ $=\left[\sum_{i=1}^{w}\left(\max\left\{-\frac{\lambda_{i}^{\ell}}{\sigma^{\ell}},g_{i}\left(\boldsymbol{x}^{\ell+1}\right)\right\}\right)^{2}\right]^{1/2}\leq\delta$ (14c)

### v2

@@ II-B Quantum Computing
This section provides the QC background and the conversion of classical QUBO problems into QC-ready forms. Then we introduce QA based on Dwave QPUs to solve QUBO problems as Fig. 1 showing, and discuss why QA is chosen over QAOA for solving SUC master problem (2).
@@ II-B1 QUBO Problem and Hamiltonian Representation
QUBO, an NP-hard problem, is one of the most widely studied optimization problems in QC. In this model, the solution to the optimization problem is composed of a series of binary variables whose value is 0 or $1$, and these variables correspond exactly to the off or on state of units in (2). The general expression of the QUBO problem is given as
@@
where $\aleph=\{1,\ldots,\mathfrak{n}\}$ denote the index set of binary variables. The variables include $\tau_{i},\tau_{j}\in\{0,1\}$, and $B_{ij},c_{i}\in\mathbb{R}$ are the coefficients of quadratic and linear terms, respectively. To enable the execution of QUBO problems on quantum hardware, it must be transformed into Ising-Lenz model. In quantum mechanics, the binary variables represent particle spin directions, where spin-up $\left|\uparrow\right\rangle=\begin{bmatrix}1&0\end{bmatrix}^{\mathrm{T}}$ and spin-down $\left|\downarrow\right\rangle=\begin{bmatrix}0&1\end{bmatrix}^{\mathrm{T}}$ correspond to $\upsilon=+1$ and $\upsilon=-1$, respectively [31]. These states are conventionally mapped to $\left|0\right\rangle$ and $\left|1\right\rangle$ in QC, leading to the variable substitution $\tau=(1-\upsilon)/2$. Based on this relation, the Ising model can be derived from (4) as
@@
While the Ising model encodes the cost function, quantum state evolution in QA and QAOA is governed by the Hamiltonian. Hence, as shown by Fig. 1, constructing a suitable Hamiltonian representation from Ising model is essential. This is achieved by mapping spin eigenstates $\upsilon$ to their eigenvalues using the Pauli-$z$ operator $\bm{\sigma}^{z}$, allowing the Hamiltonian form of the QUBO problem to be obtained by substitution in (5). Based on (5) and (6), we can construct the graphs of the QUBO problem.
$\bm{H}_{QUBO}$ $=\sum_{i\in\aleph}\sum_{j\in\aleph,i\neq j}Q_{ij}\bm{\sigma}_{i}^{z}\otimes\bm{\sigma}_{j}^{z}+\sum_{i\in\aleph}p_{i}\bm{\sigma}_{i}^{z},\;\text{with}$ (6)
$\bm{\sigma}^{z}$ $=\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right]$
II-B2 Quantum Annealer
QA is a quantum algorithm designed for quantum annealers, which leverages quantum tunneling, induced by quantum fluctuations, to escape suboptimal local minima and search for the global minimum of an objective function across a vast solution space [34]. By enabling direct transitions through energy barriers, QA can efficiently reach the global optimum, often outperforming classical algorithms in solving QUBO problems.
@@ II-B2 Quantum Annealer
The design of QA is based on the principle of adiabatic evolution. This principle states that if the system is initialized in the ground state of a known Hamiltonian, and the evolution is sufficiently slow with no energy level crossings, the system will remain in its instantaneous ground state throughout [29, 1]. When the final Hamiltonian encodes the QUBO objective function via the Ising model, the resulting ground state corresponds to the optimal solution. However, if the energy barrier is too wide, the system may still evolve to a local minimum. The mathematical expression of this process is as
$\bm{H}_{mixed}(t)=[1-s(t)]\bm{H}_{I}+s(t)\bm{H}_{F}$ (7)
@@
where $\bm{H}_{I}$ is an initial Hamiltonian, $\bm{H}_{F}$ is a final Hamiltonian that equals the objective Hamiltonian, and $s(t)$ is referred to as the normalized evolution time.
Overall, once the objective is formulated as a QUBO problem, a minor embedding is applied to map the QUBO graph onto the QPU’s sparse hardware connectivity by chaining multiple physical qubits per logical variable, and then the QPU is annealing to the lowest-energy eigenstate. The described QA procedure (7) is currently supported by D-Wave quantum annealer as illustrated in Fig. 1. In D-Wave 2000Q, qubits are implemented as tiny metal loops arranged in a Chimera lattice, where each unit cell contains 8 interconnected qubits.
Figure 1: The framework of QA using Dwave QPU to solve QUBO problems.
II-B3 Digital Quantum Computers
Digital (gate-based) quantum computers provide a general-purpose quantum computing platform based on quantum gates and circuit-based operations. A variety of quantum algorithms can be implemented on such platforms for combinatorial optimization problems.
[…]
After initialization, a depth-$d$ parameterized ansatz is constructed using alternating unitaries $\bm{U}_{I}^{d}$ and $\bm{U}_{F}^{d}$, derived from $\bm{H}_{I}$ and $\bm{H}_{F}$, respectively. These unitaries are sequentially applied and parameterized by variational angles $\bm{\beta}=(\beta_{1},\ldots,\beta_{d})$ and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{d})$, to simulate a discretized adiabatic evolution. The resulting quantum state is given in (9),
$\left|\phi(\bm{\beta},\bm{\gamma})\right\rangle=\prod_{d=1}^{\mathcal{D}}\bm{U}_{I}^{d}\bm{U}_{F}^{d}\left|\phi_{0}\right\rangle=\prod_{d=1}^{\mathcal{D}}e^{-i\beta_{d}\bm{H}_{I}}e^{-i\gamma_{d}\bm{H}_{F}}\left|\phi_{0}\right\rangle$ (9)
Finally, the parameters are updated through classical optimization to minimize the expectation value of the objective Hamiltonian, $\langle\phi(\bm{\beta},\bm{\gamma})|\bm{H}_{F}|\phi(\bm{\beta},\bm{\gamma})\rangle$, which is estimated by measuring the evolved quantum state.
II-B4 Quantum Annealers vs Digital Quantum Computers
From a practical perspective, existing studies [21, 36] indicate that the scalability of digital quantum computers for large-scale integer programming remains constrained by current hardware. First, gate-based QPU typically offer limited qubits compared to annealing-based systems. For example, the IBM Heron r2 processor provides 156 qubits, which significantly limits its scalability. Second, although QAOA can achieve asymptotic convergence of the expectation value to the ground-state energy $H_{\mathrm{min}}$ as the circuit depth increases ($\lim_{d\to\infty}\left\langle\phi(\bm{\beta},\bm{\gamma})|\bm{H}_{F}|\phi(\bm{\beta},\bm{\gamma})\right\rangle\!=\!H_{\mathrm{min}}$ [9]), practical depths are limited by noise-induced barren plateaus, where the expected gradient decays exponentially with depth and system size [47], leading to potentially suboptimal solutions. Within the BD framework, this will result in weaker lower-bound tightening such that more iterations for convergence are required.
Therefore, although gate-based approaches such as QAOA remain promising, this study focuses on near-term feasibility under current hardware constraints. Therefore, QA is adopted as the backend in this work and a simulation comparison between QA and QAOA is also provided in Appendix -A.
@@ III Proposed QPHR-ADMM Algorithm
In practice, although D-Wave QA-QPUs such as Advantage_system 4.1 provide a large qubit capacity (up to 5,627 qubits) [38], their usage remains limited, due to two main factors. (i) In minor embedding, the sparse connectivity of QA-QPU topology requires additional qubits for embedding QUBO problems, and (ii) The chain-breaking issue [37], which compromises the integrity of the QUBO representation, can lead to incorrect solutions. These challenges hinder the scalability and effectiveness of QA in large-scale optimization.
As a result, designing a QA algorithm with minimal qubit overhead is crucial for addressing real-world problems. This section first introduces binary encoding and augmented Lagrangian function to transform (2) into QUBO form and then demonstrates that a direct implementation of QA for solving (2) inevitably requires an increasing number of additional qubits as the Benders’ iterations progress, compounding the already significant qubit demand introduced by SUC itself. Two novel algorithms, namely PHR Augmented Lagrangian and quantum-based ADMM, are proposed to resolve the challenges with theoretical analysis on their computational complexity.
III-A Pre-processing for Quantization
As the master problem (2) in Section II-A is a constrained MIP problem rather than in a QUBO form, QA-QPUs cannot directly construct the corresponding quantum embedding. Therefore, preprocessing techniques are presented.
@@ III-A1 Binary Encoding
To start, the continuous variable $\Upsilon$, introduced by Benders LB (2), cannot be directly processed by the QPU. Binary encoding [53], expressed as $\Upsilon^{k}=\chi\sum_{j=0}^{J-1}2^{j}u_{j}^{k}$, can be applied where $u_{j}^{k}$ are binary variables. By selecting the appropriate precision coefficients $\chi$ and precision levels $J$, the desired precision can be attained. Leveraging the $NT$ number of binary variables in the original SUC, the qubit occupation after binary encoding becomes $\underline{NT+J}$.
@@ III-A2 Augmented Lagrangian Function
In the UC formulation and the Benders decomposition framework, inequality constraints are typically handled by incorporating them into the objective function through augmented Lagrangian techniques, which will introduce extra binary encoded slack variables for execution on a QPU. This applies both to structural UC constraints, such as minimum up/down time constraints, and to the Benders cuts generated during the iterative process.
Specifically, suppose there are $\mathcal{M}$ minimum up/down time constraints, each requiring $\mathcal{P}$ qubits for binary-encoded slack variables to achieve the desired precision. This results in a fixed qubit overhead of $\mathcal{M}\mathcal{P}$. In addition, according to (2c), the number of Benders cuts grows with the iteration index $k$. Assuming the precision level of each corresponding slack variable is $\mathcal{F}$, the resulting qubit overhead increases linearly with $k$ and can be expressed as $\underline{NT+J+\mathcal{M}\mathcal{P}+k\mathcal{F}}$.
Considering the limited number of qubits in quantum computers, introducing slack variables is inefficient and unacceptable. This inefficiency serves as the primary motivation for the PHR Lagrangian algorithm proposed in this Section III-B. Moreover, the qubit occupation grows linearly with the number of generator $N$. This qubit overhead will be resolved by the quantum ADMM algorithm in Section III-C.
@@ III-B1 Classic Formulation
The Powell-Hestenes-Rockafellar Augmented Lagrangian Multiplier (PHR-ALM) method was originally designed for equality-constrained optimization problems. Rockfellar later extended it to handle inequality constraints by eliminating the need for slack variables in optimization problems [40], which is an ideal property for QA.
@@
$\min$ $f(\bm{x})$ (10)
s.t. $g_{i}(\bm{x})\leq 0\text{ for }i=1,\ldots,w$
where $n$ represents the total number of unit states and binary-encoded variables. $\bm{x}\in\{0,1\}^{n}$ and $w$ is the total number of binary inequality constraints after some BD iterations. After introducing slack variables $\mathcal{S}_{i}$ to relax each inequality constraints, the augmented Lagrangian function $\mathcal{L}$ of (10) can be given as
$\!\mathcal{L}$ $=f(\bm{x})\!+\kappa(\bm{x},\mathcal{S},\lambda,\sigma)$ (11)
$=f(\bm{x})\!+\!\sum_{i=1}^{w}\lambda_{i}\!\left[g_{i}(\bm{x})\!+\mathcal{S}_{i}{}^{2}\right]\!+\frac{\sigma}{2}\sum_{i=1}^{w}\left[g_{i}(\bm{x})\!+\mathcal{S}_{i}{}^{2}\right]^{2}$
@@
$\frac{\partial\mathcal{L}}{\partial\mathcal{S}_{i}}=2\lambda_{i}\mathcal{S}_{i}+2\sigma\mathcal{S}_{i}\left[g_{i}(\bm{x})+\mathcal{S}_{i}^{2}\right]=0,\quad\forall i.$ (12)
@@
${\mathcal{S}_{i}{}^{2}}=\begin{cases}\hfil{-\frac{\lambda_{i}}{\sigma}-g_{i}(\bm{x})},&\lambda_{i}+\sigma g_{i}(\bm{x})\leq 0\\ \hfil{0,}&{\lambda_{i}+\sigma g_{i}(\bm{x})>0}\end{cases}$ (13)
From (13), when $\lambda_{i}+\sigma g_{i}(\bm{x})\leq 0$, $\mathcal{S}_{i}{}^{2}$ is non-negative, which means that the solution satisfies the $i$-th constraint $g_{i}(\bm{x})\leq 0$. Otherwise, it indicates that the solution is not within the feasible region. Next, substituting the above results into the augmented penalty function, the corresponding PHR-based expression is given as follows,
${\kappa_{i}}=\begin{cases}\hfil-\frac{\lambda_{i}{}^{2}}{2\sigma},&\lambda_{i}+\sigma g_{i}(\bm{x})\leq 0\\ \hfil{\frac{1}{2\sigma}\left[\left(\sigma g_{i}(\bm{x})+\lambda_{i}\right)^{2}-\lambda_{i}{}^{2}\right],}&{\lambda_{i}+\sigma g_{i}(\bm{x})>0}\end{cases}$ (14)
@@
[…]
@@
At iteration $\ell$, solving the unconstrained PHR augmented Lagrangian function (15) yields the solution $\bm{x}^{\ell}$. Let $\mathcal{R}$ denote the constraint residual. If $\bm{x}^{\ell}$ satisfies the stop criterion (16c), the iteration is terminated, and the output $\bm{x}^{\ell}$ is the approximate minimum solution of the original problem. Otherwise, the Lagrange multipliers $\lambda_{i}^{\ell+1}$ and the penalty parameter $\sigma^{\ell+1}$ are updated according to (16a) and (16b). Here, $\varrho\in(0,1)$ is a residual reduction factor that determines whether sufficient progress in the residual is achieved, and $\eta>1$ is the penalty growth factor.
$\lambda_{i}^{\ell+1}$ $=\max\!\left\{\lambda_{i}^{\ell}+\sigma^{\ell}g_{i}\left(\bm{x}^{\ell+1}\right),0\right\},\quad i=1,\ldots,w$ (16a)
$\sigma^{\ell+1}$ $=\begin{cases}\sigma^{\ell},&\text{if }\mathcal{R}^{\ell+1}<\varrho\mathcal{R}^{\ell},\\[2.0pt] \eta\,\sigma^{\ell},&\text{otherwise}\end{cases}$ (16b)
$\mathcal{R}^{\ell}$ $=\left[\sum_{i=1}^{w}\left(\max\left\{-\dfrac{\lambda_{i}^{\ell}}{\sigma^{\ell}},g_{i}\left(\bm{x}^{\ell+1}\right)\right\}\right)^{2}\right]^{1/2}\leq\delta$ (16c)